If a crew of 3 people take 2 ½ hours to clean a house, then how long should a crew of 2 take to clean the same house? Assume that all people in the cleaning crew work at the same steady rate.

Answer by stanbon(75887)

If a crew of 3 people take 2 ½ hours to clean a house, then how long should a crew of 2 take to clean the same house? Assume that all people in the cleaning crew work at the same steady rate.
----
# of hrs and # of people are indirectly ralated.
----
hrs = k/p
Solve for "k" using "If a crew of 3 people take 2 ½ hours to clean a house"
---
(5/2) = k/3
k = 15/2
---
Equation:
hrs = (15/2)/p
---
how long should a crew of 2 take to clean the same house?
hrs = (15/2)/2
hrs = 15/4 = 3 3/4 hrs
============================
Cheers,
Stan H.
==================